# Symmetry-Projected Spin-AGP Methods Applied to Spin Systems

### v1

@@ IV Conclusions
In this article, we have studied symmetry-projected AGP methods for the 1D XXZ model and the 2D $\mathrm{J_{1}-J_{2}}$ model with square and triangular lattices. The projection of space group symmetry is achieved with a linear combination of permuted AGPs, resulting in a computational cost of $\mathcal{O}(M^{4})$, which is one factor of $M$ more expensive than a single mean field optimized AGP. The symmetry-projected AGP is optimized with the variation after projection approach. It can be seen that the space group projected AGP not only restores the correct symmetry of the system but adds important strong correlations that a single optimized AGP misses.
@@ Acknowledgements.
This work was supported by the U.S. Department of Energy, Office of Basic Energy Sciences, Computational and Theoretical Chemistry Program under Award DE-SC0001474. G.E.S. acknowledges support as a Welch Foundation Chair (Grant No. C-0036).

### v2

@@ IV Conclusions
In this article, we have studied symmetry-projected AGP methods for the 1D XXZ model and the 2D $\mathrm{J_{1}-J_{2}}$ model with square and triangular lattices. The projection of space group symmetry is achieved with a linear combination of permuted AGPs, resulting in a computational cost of $\mathcal{O}(M^{4})$, which is one factor of $M$ more expensive than a single mean field optimized AGP. The symmetry-projected AGP is optimized with the variation after projection approach. It can be seen that the space group projected AGP not only restores the correct symmetry of the system but also adds important strong correlations that a single optimized AGP misses.
@@ Acknowledgements.
This work was supported by the U.S. Department of Energy, Office of Basic Energy Sciences, Computational and Theoretical Chemistry Program under Award DE-SC0001474. G.E.S. acknowledges support as a Welch Foundation Chair (Grant No. C-0036). We thank Dr. Ruiheng Song for generously providing the sBCS-CCSD results shown in Fig. 6.